DAY 16: 17/07/17
After a 2 weeks break, where I went to London to visit my friend, I'm back in the lab, and excited for the next part of my internship project.
Today I did miniprep DNA for colony PCR using N1, N3, N5, N8 and HBT plasmids.
INFO: what was previously called plasmids 1, 3, 5, 8, 10 and b3.4 are now called N1, N3, N5, N8, N10 and HBT, respectively.
The PCR was prepared in the following order: MW maker, N1, N3, N8, N10, N5 and HBT, where NSEQ1-RS1 primer mix was used for N1, N3, N8 and N10 DNAs; FS1-RS2 primer mix used for N5 DNA; and 5NSEQ-HBT-PCR-REV primer mix used for HBT DNA.
DAY 17: 18/07/17
The PCR prepared yesterday was ran on agarose gel to see the results. The results are shown in figure 10.
Figure 10: Agarose gel analysis of DNA MW marker against PCR solutions N1, N3, N8, N10, N5 and HBT, respectively
As it can be seen from figure 10, the PCR didn't really work for most of the DNA samples, since only the primers are giving out a signal for DNA samples N1, N3, N8, N10. However, for DNA sample N5, there are two visible strands, suggesting that PCR has fully worked, showing the prier strand and the DNA part of interest strand.
To check whether the DNA samples were ok, a new PCR was prepared. This time, using FS1-RS2 and NSEQ1-RS1 primer mixes. Results are shown below in figure 11.
Figure 11: Agarose gel analysis of DNA MW marker against PCR solutions N1, N3, N8, N10, N5, HBT (treated with FS1-RS2 primer mix), N1, N3, N8, N10, N5, HBT (treated with NSEQ1-RS1 primer mix), respectively.
As it can be seen from figure 11, the DNA samples are ok because, at least for FS1-RS2 primers, all DNA samples gave signal at the point of interest, suggesting the primers were the problem on the previous PCR reaction. Furthermore, it can be seen that the solutions treated with NSEQ1-RS1 primer mix didn't give signal for the DNA area of interest.
To check whether it was the primer not working, a new primer mix with NSEQ1-RS1 primers was prepared to do a new PCR.
DAY 18: 19/07/17
The results from yesterday's PCR is shown in figure 12.
It can be seen from figure 12 that the PCR didn't work, as only the primers themselves are emitting signal, thus, suggesting that the NSEQ1-RS1 primers combination is not binding properly to the DNA's part of interest.
In order to identify which primer is not working properly, new primer combinations were prepared and tested on N1 and N3 DNAs. The first new primer mix combination is NSEQ1-RS2 (RS2 have worked before with FS1 primer, as shown in figure 10 and 11). The second new primer mix combination is 5NSEQ-RS1. Results are shown in figure 13.
From figure 13, it can be seen that N1 DNA treated with NSEQ1-RS2 primers, show a very light DNA band. This suggests that some DNA from the part of interest is giving out signals. Furthermore, N3 DNA treated with NSEQ1-RS2 primers did not give any signal other than the primer itself. Neither N1 or N3 DNAs, treated with 5NSEQ-RS1 primers, gave any signals. These results suggest that the NSEQ1 primer is the problem in the NSEQ1-RS2 primer mix and we don't have enough data to determine which primer is the problem from the 5NSEQ-RS1 primer mix (results from figure 10, regarding the HBT treated with 5NSEQ/HBT-PCR-REV were inconclusive).
In other to obtain more data, other colonies of the HBT tag colonies were treated with 5NSEQ-HBT0PCR-REV primer mix. Results are shown in figure 14.
INFO: The primer mix should only bind if the HBT tag is present in the DNA.
Figure 14: Agarose gel analysis of DNA MW marker against HBT tag colonies 3.1, 3.2, 3.4, 3.5, 3.6, treated with 5NSEQ-HBT-PCR-REV primer mix.
It can be seen from figure 14 that out of 5 HBT tag colonies, only one seems to actually have the HBT tag. The colony with a HBT tag seems to be colony #5-2.
DAY 19: 20/07/17
The only colony with a HBT tag seems to be colony #5-2, in figure 14, but since none of the others appear to have the HBT tag, which they should have (in theory), a new PCR with all #5 colonies and #6 colonies, as a control, was ran with different primers. The colonies were treated with NRP1-K185-REV primer mix, and also NRP3-NRP4 primer mix. There are two colonies for #5 and two for #6, so they were labelled #5-1, #5-2, #6-3, #6-4, and each colony was treated with each set of primer mix.
Furthermore, from figure 10, it can be seen that only N5 had a DNA signal, so DNA from all N5 and N6 colonies were prepared and treated with FS1-RS2 primer mix for PCR. There are four colonies within N5 and four within N6, so they were labeled N5-1, N5-2, N5-3, N5-4, N5-6, N6-6, N6-7 and N6-8.
The results for all #5, #6, N5 and N6 colonies are shown in figure 15.
From figure 15 results, it can be seen that the NRP1-K185-REV primer mix didn't work, as there is no DNA signal visible. However, that was expected because primer K185-REV optimal annealing temperature is lower than 60oC, so a new PCR with annealing temperature at 48oC was ran for #5-1, #5-2, #6-3 and #6-4 DNAs treated with NRP1-L185-REV.
DAY 20: 24/07/17 The results from DAY 19's PCR is shown in figure 16. It can be seen from figu...
DAY 1: 6/6/17 I was super nervous for my first day and I didn't know what to really expect! Once I arrived, I was greeted by Dr. MacNe...
This summer, I have secured an internship for the Summer of 2017 in the School of Biology at the University of St Andrews, under the supe...
DAY 5: 12/06/17 Today's plan was to confirm whether plasmids 3 and 7, in fact, contained the insert at the correct orientation. This wi...